OpenStudy (anonymous):

Could someone please help me with this? :) log5x + log4 = 3 HINT: Notice that the logs are in base 10.

OpenStudy (mathstudent55):

There is a rule of logs that is helpful here. \( \log a + \log b = \log ab\) Can you apply that rule to the left side?

OpenStudy (ankit042):

log 20x = 3

OpenStudy (mathstudent55):

Once you apply that rule, then change the log equation into the exponential form using this: \(\log_b x = y \iff b^y = x\)

OpenStudy (anonymous):

Okay that makes sense.